A 440-V, 60-Hz, 4-pole, three-phase, A-connected, synchronous motor has
a synchronous impedance of 1.2 + j8 n/phase. The field current is adjusted
so that the excitation voltage is 560 V. When the motor develops the maximum
power, determine (a) the line current, (b) the power factor, (c) the
torque angle, (d) the maximum power developed, and (e) the maximum
torque developed.
Answer #1

RD. ELO Griven data N440 V 792 JTa Ta = ZAL 4 Po 1e 3- Pouun tlow ea ab hmotey P= ReL EI] Cop (8-6)-Cop Zg=L2+ 8J/Phas 28= 8.b) A2 we know that Angliblw IVlaI, Prwn Bgle Anple b/w Elane 1 Ia whane S7-126-+61 469 -2 4.271 - Ve sign Lending ghaws P.f 212-0 Xt NS P-TXW PmakTimar X 24 709 KW Tm AX e) 120x 60 4 Tm Ax 24 719XID 6 0 2 4.709X D X6D 2 XI6D Ans Nm -Tmax 131 090 60 1hope this will help